• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 07:58:49 on April19,2007

    Entry number 199568

    keyword=Owl shift summary

    No beam at beginning of shift.

    At about 1:40 we get the beam back. We started by getting the beam position to spec, then it took some time to get the raster size properly adjusted, because there was some confusion in the fact that MCC's "width" and "height" corresponded to our Y and X, respectively.

    At about 3:15 we finished adjusting the raster size and moved to 4He, and we started ramping up the current. After a little while we maxed out at about 73 muA, limited by the CHL flow and keeping the HPH heat load on the target well above zero.

    We have nearly completed the program on the white board. At present we are running on 4He with the 6 msr collimator on the left HRS.